A probability experiment consists of rolling an 8​-sided die. Find the probability of the event below. rolling a number with exactly 3 not necessarily unique prime factors The probability is . 38.
Answer #1

A 8 sided die is rolled . The outcomes are 1,2,3,4,5,6,7,8.

4= 1*2*2

6= 1*2*3

8= 1*2*4

Only 4,6 and 8 can be represent in exactly 3 not necessarily unique prime factors.

Therefore, Probability that rolling a number with exactly 3 not necessarily unique prime factors = 3/8 = 0.375=0.38( approximately).